Given the following data with n=5, compute the covariance sxy. Give an exact answer. x 2 3 5 9 11

y -21 -15 -18 -10 -5

Note that ∑i=15xi=30, ∑i=15x2i=240, ,∑i=15yi=−69, ∑i=15y2i=1115
Answer #1

the necessary calculation table:-

xi yi x_{i}y_{i}
2 -21 -42
3 -15 -45
5 -18 -90
9 -10 -90
11 -5 -55
sum= -322

the covariance (sample) is :-

677-6( “ = fis

=\frac{1}{5-1}[-322-\frac{30*-69}{5}]

=\mathbf{23}

i am calculating the population covariance also,for your better understanding.

\sigma_{xy}=\frac{1}{n}[\sum x_{i}y_{i}-\frac{\sum x_{i}\sum y_{i}}{n}]

  30 * -69,

  =\mathbf{18.4} (this is the population covariance)

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible...